OpenStudy (anonymous):

Graph y=-4x^2-2 And Its Inverse

hero (hero):

To find the inverse of y=-4x^2-2, Swap x and y, then isolate y: x = -4y^2 - 2 4y^2 = -x - 2 y^2 = -(x + 2)/4 y = + or - sqrt(-(x + 2))/2

OpenStudy (anonymous):

Thank You c:

hero (hero):

yw